# OnePerc: A Randomness-aware Compiler for Photonic Quantum Computing

### v2

@@ Abstract.
The photonic platform holds great promise for quantum computing. Nevertheless, the intrinsic probabilistic characteristics of its native fusion operations introduces substantial randomness into the computing process, posing significant challenges to achieving scalability and efficiency in program execution. In this paper, we introduce a randomness-aware compilation framework designed to concurrently achieve scalability and efficiency. Our approach leverages an innovative combination of offline and online optimization passes, with a novel intermediate representation serving as a crucial bridge between them. Through a comprehensive evaluation, we demonstrate that this framework significantly outperforms the most efficient baseline compiler in a scalable manner, opening up new possibilities for realizing scalable photonic quantum computing.
@@ 1. Introduction
Photonic platform holds great promise for universal quantum computing due to the unique advantages of photonic qubits (OBrien2009, ; Bogdanov:17, ), including their great scalability, long coherence time and easy integration with quantum networks. In addition to the experimental demonstration of quantum supremacy on photonic systems (zhong2020quantum, ; PhysRevLett.127.180502, ; Madsen2022, ), Xanadu (xanadu, ) has released their programmable nanophotonic chips that can serve as a new testbed, while PsiQuantum (fbqc, ) has proposed their technology roadmap towards one million qubits using silicon photonics.
@@
Photonic quantum computing differs from other platforms such as superconducting (devoret2013superconducting, ), ion trap (bruzewicz2019trapped, ) and neutral atoms (saffman2016quantum, ), as it is scaled up by a probabilistic operation known as fusion (fbqc, ). This probabilistic feature comes intrinsically from the degeneracy in fusions’ outputs for different input states, bringing significant randomness to the computing process. On the hardware side, improvement of fusion success probability to a high value requires an impractical amount of ancillary resources (0.78, ; 0.95with30, ). On the software side, this randomness is not taken into consideration by existing software infrastructures for the circuit-based programming model  (nielsen2001quantum, ), (e.g., Qiskit (Qiskit, ), Tket (tket, )), as photonic systems adopt a different computing model, which is known as measurement-based quantum computation (MBQC) or one-way quantum computation (1WQC) (mbqc2009, ; translation, ).
@@
Recently, as an initial effort towards efficient photonic MBQC, a compilation framework OneQ (OneQ, ) has been proposed to significantly reduce the number of required fusions and the overall program execution time. However, it overlooks the severe randomness introduced by fusion failures, simply assuming that fusions always succeed. As a compiler, OneQ translates the construction of a program-specific entangled state called graph state  (translation, ) (Fig. 1(a)) into a fusion pattern between the small entangled resource states (fbqc, ) available on the hardware (Fig. 1(b)). However, when fusion failures occur in real-time execution, as illustrated in Fig. 1(b), the resulting state becomes a random graph state deviating from the target structure. Thus the execution needs to be retried untill success, which is non-scalable given that a practical fusion success probability in the near term is merely around 75% (0.78, ; 0.95with30, ). From now on, we will refer to the graph states required by programs as program graph states and those generated by fusions as physical graph states.
Figure 1. Randomness brought by fusion failures on a photonic quantum hardware with 3-qubit resource states.
@@
The objective of this paper is to present a scalable and efficient compilation framework capable of effectively handling fusion failures in the computing process. This is intuitively a hard problem. Firstly, with a failure rate as high as 25%, it seems impossible to ensure the formation any specific entanglement structure among the photonic qubits. Secondly, failed entanglements such as the disconnected edges in Fig. 1 cannot be recovered since the photons involved in a fusion are completely destroyed by the fusion. Thirdly, the limited lifetime of photons refuses the execution of over-complex algorithms in real-time, as photons are prone to an increasing loss rate with prolonged storage time in fibers (fiber, ).
@@
Fortunately, there are some nice features of fusions and graph states that we can leverage. Firstly, fusion failures are heralded (linear_optics, ), allowing real-time awareness and enabling the incorporation of classical feed-forward (linear_optics, ; 99feedforward, ; nonlinear_feedforward, ) to adjust subsequent operations based on prior fusion outcomes. Secondly, when the fusion success probability exceeds a threshold, the resulting physical graph state contains a long-range-connected component with a high probability. This widely studied phenomena, known as percolation (3GHZ, ; perc_threshold, ; 2D_renorm, ), plays a crucial role in providing viable computing resource, inspiring our framework’s name, OnePerc (one-way quantum computing based on percolation). Thirdly, a random graph state can be reshaped into any subgraph of it by eliminating the redundant qubits, which can be achieved by measuring them out in $Z$-bases (mbqc2009, ).
@@
However, leveraging these features is highly non-trivial. In addition to the absence of a generic fusion strategy among general resource states for achieving percolation and the structural mismatch between the high-level program graph state and the low-level random physical graph state, we need to always keep in mind the limited time for real-time passes. Specifically, the process associated with the formation of long-range connectivity and the reshaping of the random physical graph state both need to be carried out in real-time, leading to a high demand on their lightweight design. This results in a conflict between the real-time scalability and the program execution efficiency, as a sophisticated optimization for program execution requires complex algorithms that are not feasible in real-time.
@@
To this end, we propose a randomness-aware compiler designed to efficiently scale up quantum computing on photonic systems, as illustrated in Fig. 2. Our framework achieves concurrent real-time scalability and program execution efficiency through the combination of an online pass and an offline pass. The online pass handles real-time randomness in a scalable manner through percolation and reshaping. In particular, it provides a generic fusion strategy for various resource states and mitigates the structural mismatch between program and physical graph states by enabling the abstraction of a virtual hardware. Motivated by the features of this virtual hardware, we propose a FlexLattice intermediate representation (IR), which preserves the high-level program information and provides maximal optimization space supported by the hardware. This allows offline passes to optimize the program execution by transforming the program to an efficient IR program, with the IR program translated to a low-level instruction set to guide real-time operations.
@@
  • We propose a randomness-aware compilation framework for photonic quantum computing through an innovative combination of online and offline passes, which are bridged by a novel FlexLattice IR facilitated with a low-level instruction set.
  • The online pass handles real-time randomness in a scalable manner by enabling long-range connectivity formation among general resource states and efficient reshaping of the random long-range connected computing resource.
  • The FlexLattice IR provides programs with a maximal optimization space supported by the underlying hardware, allowing for the introduction of a sophisticated offline pass to improve the efficiency of program execution.
@@
2. Background and Motivation
@@
2.1. Photonic MBQC
Photonic systems are particularly suitable for MBQC (con-rev, ), a conceptually distinct computating paradigm from the circuit model. In MBQC, computation is driven by single-qubit projective measurements, rather than 1-qubit and 2-qubit gates, on an initial entangled state called graph state, whose graph structure $G=(V,E)$ is determined by the quantum program (translation, ), as exemplified in 3(a). Specifically, a graph state consists of qubits located on $G$, formally defined as the eigenstates of operator
$s=X_{i}\bigotimes_{j\in n_{i}}Z_{j},\quad\forall i\in V$
where $n_{i}$ is the set of neighboring qubits of $i\in V$. The measurement basis of each qubit is predetermined by the quantum program, known as a measurement pattern, but are subject to a real-time adjustment according to the measurement outcomes of prior qubits. This feed-forward is used to address the non-determinism of measurement outcomes.
The feasibility of photonic MBQC has been demonstrated with various quantum algorithms on photonic qubits. For instance, Grover’s algorithm achieved 0.9 fidelity on a 4-photon 4-qubit graph state (mbqc-grover, ), while the Deutsch-Jozsa algorithm reached 0.96 fidelity on a 6-qubit graph state (mbqc-dj, ). In addition, Simon’s algorithm was successfully applied to a five-qubit cluster state for period-finding(mbqc-simon, ). Notably, the photonic platform is rapidly scaling up with integrated waveguides and optical chips  (con-rev-217, ; con-rev-204, ; ferreira2022deterministic, ; con-rev-247, ; con-rev-252, ), enabling the realization of a programmable optical processor (xanadu, ).
2.2. Realistic Hardware
Practical photonic hardware scales up by creating small resource states, e.g., 4-qubit, 6-qubit graph states, and connecting them through fusions (fbqc, ). In particular, identical resource states are periodically generated by an array of resource state generators (RSG) every cycle, with those generated in the same RSG cycle forming a resource state layer (RSL), as depicted by Fig. 3(b). These resource states are merged probabilistically into larger graph states through (type II (linear_optics, )) fusions, which are concurrent measurements of $X\otimes Z$ and $Z\otimes X$ on two photonic qubits from different resource states. Resource states on the same RSL can fuse with their neighbors, as shown in Fig. 3(c), by a spatial routing of photons, while those on different RSLs generated by the same RSG can fuse with each other by a temporal routing that controls photons’ arrival times at measurement devices.
With the advanced integrated silicon photonics, hardware components described above can operate on the scales of GHz clock rates (GHz_1, ; GHz_2, ; GHz_3, ), leading to a time scale $\sim$ 1 ns for RSG cycles. Spatial routing can be adjusted in every RSG cycle with switches, while temporal routing can be achieved by temporarily storing photonic qubits in a high-capacity quantum memory known as delay lines, which can be realized with current optical fiber technology. With a low transmission loss rate of $<5\%$ per km (fiber, ), photons can have a lifetime of around 5000 RSG cycles in the delay lines. Moreover, the size of RSL is not completely constrained by the number of RSGs, but can be extended by leveraging the tradeoff between spatial and temporal fusions (interleaving, ), with the 5000-cycle photon lifetime allowing a 5000$\times$ extension at most.
However, as the key operation for merging resource states, fusions are intrinsically probabilistic. By allowing two ancilla photons, their success probability can be practically boosted to 75% (0.78, ; 0.95with30, ). While no conceptual limit has been found yet, so far the maximum known success probability attainable using linear optics is 78% with 8 ancilla single photons (0.78, ). Reaching a higher success probability not only requires a larger number of ancilla photons but also require the ancilla photons to be entangled. For example, it would take 30 ancilla entangled photons to reach a probability over 95% (0.95with30, ).
2.3. Motivating Example
Previous work has proposed OneQ (OneQ, ), an efficient compilation framework for photonic MBQC. We will show that a straightforward adaption of OneQ is insufficient to yield a scalable compiler in the presence of fusion failures.
@@ 2.3. Motivating Example
Fig. 3(c) illustrates OneQ’s strategy with an example of generating a graph state in Fig. 3(a) from a single RSL depicted in Fig. 3(b), with the strategy represented more compactly by Fig. 3(d). To handle real-time randomness, a straightforward adaption is to introduce a retry mechanism. For example, the strategy in Fig. 3(c) can result in a dynamic implementation in Fig. 3(e) according to the fusion successes and failures (green and red ellipses), with these fusions performed sequentially, from t1 to t11. If a fusion, such as t3, fails, we retry the fusion using another two qubits at t4, and the same approach is applied to t6 and t7. This allows us to successfully generate the graph state in Fig. 3(a).
However, it is worth noting that some fatal failures may necessitate the retrial of the entire compilation. For example, in Fig. 3(f), the triangular structure ABC is successfully generated from t1 to t4, but subsequent failures at t5 and t6 deplete the qubits in ABC, only leaving the isolated qubit B. In Fig. 3(g), a 5-qubit linear graph state forms from t1 to t11, which provides the potential for generating the triangle ABC if the fusion at t11 succeeds in fusing the two qubits at the line ends. Unfortunately, this fusion fails and consumes the last neighboring resource state of qubit A (except C), leaving A with no chance to fuse with other qubits.
@@
From this example, we can find some critical issues of this dynamic retry mechanism. First, adapting to prior fusion outcomes necessitates a sequential execution of fusions. This considerably extends the processing time for each resource state layer (RSL), resulting in a time inefficiency as subsequent RSLs must wait for the completion of the current one. Second, since the decision-making process for responding to prior fusions occurs in real time, this extended processing time could exceed the limited lifetime of photons, especially for large RSLs. This would result in substantial photon loss, compromising the overall fidelity as computing scales up. Third, the frequent retries in real-time implementation lead to significant deviations from the planned strategy in Fig. 3(c). This undermines the benefits of the proactive planning, eroding the efficiency achieved by the mapping strategy of OneQ.
3. Overview
Tolerating randomness in the compilation while maintaining efficiency presents a significant challenge. To address this, we propose an innovative framework that achieves scalability and efficiency simultaneously through a synergy of online and offline passes. The online pass prioritizes the real-time scalability by maximizing the concurrency among fusions and the parallelism of the algorithm for decision making. The offline pass focuses on the efficient deployment of high-level program graph states onto the randomness-eliminated computing resource guaranteed by the online pass. The bridge between the online and offline passes is established through an intermediate software layer positioned between the low-level physical layer and the high-level program layer. This motivates a novel FlexLattice IR, along with an instruction set supported by the online pass and fulfilling the requirements of the offline pass.

### v3

@@ Abstract.
The photonic platform holds great promise for quantum computing. Nevertheless, the intrinsic probabilistic characteristic of its native fusion operations introduces substantial randomness into the computing process, posing significant challenges to achieving scalability and efficiency in program execution. In this paper, we introduce a randomness-aware compilation framework designed to concurrently achieve scalability and efficiency. Our approach leverages an innovative combination of offline and online optimization passes, with a novel intermediate representation serving as a crucial bridge between them. Through a comprehensive evaluation, we demonstrate that this framework significantly outperforms the most efficient baseline compiler in a scalable manner, opening up new possibilities for realizing scalable photonic quantum computing.
journalyear: 2024copyright: rightsretainedconference: 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 3; April 27-May 1, 2024; La Jolla, CA, USAbooktitle: 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 3 (ASPLOS ’24), April 27-May 1, 2024, La Jolla, CA, USAdoi: 10.1145/3620666.3651372isbn: 979-8-4007-0386-7/24/04
@@ 1. Introduction
Photonic platform holds great promise for universal quantum computing due to the unique advantages of photonic qubits (OBrien2009, ; Bogdanov:17, ), including their great scalability, long coherence time and easy integration with quantum networks. Besides the experimental demonstration of quantum supremacy on photonic systems (zhong2020quantum, ; PhysRevLett.127.180502, ; Madsen2022, ), PsiQuantum has proposed their technology roadmap towards one million qubits using silicon photonics (fbqc, ; interleaving, ). The potential high clock speed of this approach (interleaving, ) could make photonic platform advantageous for near-term quantum algorithms.
@@
Photonic quantum computing differs from other platforms such as superconducting (devoret2013superconducting, ), ion trap (bruzewicz2019trapped, ) and neutral atoms (saffman2016quantum, ), as it is scaled up by a probabilistic operation known as fusion (fbqc, ). Fusion plays a key role of forming large-scale entanglements between photonic qubits by merging small entangled states into larger ones upon success. Its probabilistic feature comes intrinsically from the degeneracy in fusions’ outputs for different input states (0.95with30, ), bringing significant randomness to the computing process. On the hardware side, improvement of fusion success probability to a high value requires an impractical amount of ancillary resources (0.78, ; 0.95with30, ). On the software side, this randomness is not taken into consideration by existing software infrastructures for the circuit-based model  (nielsen2001quantum, ) (e.g., Qiskit (Qiskit, ), Tket (tket, )). This is because the weak interaction between photons makes it hard to realize 2-qubit gates in the circuit model, but favors a different computing model known as measurement-based quantum computation (MBQC) or one-way quantum computation (1WQC) (mbqc2009, ; translation, ).
@@
Recently, as an initial effort towards efficient photonic MBQC, a compilation framework OneQ (OneQ, ) has been proposed to significantly reduce the depth of compiled programs and the number of required fusions. However, it overlooks the severe randomness brought by fusion failures, simply assuming that fusions always succeed. As a compiler, OneQ translates the construction of a program-specific entangled state called graph state  (translation, ) (Fig. 1(a)) into a fusion pattern between the small entangled resource states (fbqc, ) available on the hardware (Fig. 1(b)). However, when fusion failures occur in real-time execution, as illustrated in Fig. 1(b), the resulting state becomes a random graph state deviating from the target structure. Thus the execution needs to be retried until success, which is non-scalable given that a practical fusion success probability in the near term is merely around 75% (0.78, ; 0.95with30, ). From now on, we will refer to the graph states required by programs (e.g., Fig. 1(a)) as program graph states and those generated by fusions (e.g., Fig. 1(c)) as physical graph states.
Figure 1. Randomness brought by fusion failures.
@@
The objective of this paper is to present a scalable and efficient compilation framework capable of effectively handling fusion failures in the computing process. This is intuitively a hard problem. Firstly, with a failure rate as high as 25%, it seems impossible to ensure the formation of any specific entanglement structure among the photonic qubits. Secondly, failed entanglements such as the disconnected edges in Fig. 1 cannot be recovered since the photons involved in the fusion are completely destroyed by the fusion. Thirdly, the limited lifetime of photons refuses the execution of over-complex algorithms in real-time, as photons are prone to an increasing loss rate with prolonged storage time in fibers (fiber, ).
@@
Fortunately, there are some nice features of fusions and graph states that we can leverage. Firstly, fusion failures are heralded (linear_optics, ), allowing real-time awareness and enabling the incorporation of classical feed-forward (linear_optics, ; 99feedforward, ; nonlinear_feedforward, ) to adjust subsequent operations based on prior fusion outcomes. Secondly, when the fusion success probability exceeds a threshold, the resulting physical graph state contains a long-range-connected component with a high probability. This widely studied phenomenon, known as percolation (3GHZ, ; perc_threshold, ; 2D_renorm, ), plays a crucial role in providing viable computing resource, inspiring our framework’s name, OnePerc (one-way quantum computing based on percolation). Thirdly, a random graph state can be reshaped into any subgraph of it by eliminating the redundant qubits, which can be achieved by measuring them out in $Z$-bases (mbqc2009, ).
@@
However, leveraging these features is highly nontrivial. In addition to the absence of a general fusion strategy to achieve percolation for generic resource states, and the structural mismatch between the high-level program graph state and the low-level random physical graph state, we need to keep in mind the limited time for real-time passes. Specifically, the process associated with the formation of long-range connectivity and the reshaping of the random physical graph state both need to be carried out in real-time, leading to a high demand on their lightweight design. This creates a conflict between the real-time scalability and the program execution efficiency, as a flexible optimization strategy for efficient program execution may require complex algorithms that are not feasible in real-time.
@@
To this end, we propose a randomness-aware compiler to efficiently scale up quantum computing on photonic systems, as illustrated in Fig. 2. Our framework achieves concurrent real-time scalability and program execution efficiency through the combination of an online pass and an offline pass. The online pass handles real-time randomness in a scalable manner through percolation and reshaping. In particular, it provides a general fusion strategy for various resource states and exposes the reshaped physical graph states to programs by the abstraction of a virtual hardware. Motivated by the features of this virtual hardware, we propose a FlexLattice intermediate representation (IR), which preserves the high-level program information and provides maximal optimization space supported by the virtual hardware. This allows offline passes to address the mismatch between program and physical graph states by transforming the program to an efficient IR program, which can then be translated to intermediate-level instructions to guide real-time operations.
@@
  • We propose a randomness-aware compiler for photonic quantum computing through a combination of online and offline passes, which are bridged by a novel FlexLattice IR facilitated with an intermediate-level instruction set.
  • The online pass handles real-time randomness in a scalable manner by formation of long-range connectivity with various resource states and efficient reshaping of the random long-range connected structures.
  • The FlexLattice IR provides programs with an optimization space that balances the complexity of online structure reshaping and the flexibility of the reshaped structures. This enables an offline pass to enhance the efficiency of program execution through optimized mapping algorithms.
@@
2. Background
2.1. MBQC Basics
MBQC is a universal but conceptually distinct computational model from the circuit model. In MBQC, computation is driven by 1-qubit projective measurements, rather than 1-qubit and 2-qubit gates, on an initial entangled state called graph state, whose graph structure $G=(V,E)$ is determined by the quantum program (translation, ). As exemplified in 1(a), 3(b), 5(a), each vertex in the graph state stands for a qubit, with the state formally defined as the eigenstates of operator
$s=X_{i}\bigotimes_{j\in n_{i}}Z_{j},\quad\forall i\in V$
where $X_{i},Z_{j}$ are the Pauli operators on qubit $i,j$ respectively, and $n_{i}$ is the set of neighboring qubits of $i\in V$ on graph $G$. On the graph state, computation can be driven by a set of equatorial measurements $E(\alpha)$, i.e., measurements on the X-Y plane of Bloch sphere at an angle $\alpha$, along with $Z$-measurements. The measurement basis of each qubit is predetermined by the quantum program, known as a measurement pattern, but are subject to a real-time adjustment according to the measurement outcomes of prior qubits, with the angles adjusted from $\alpha$ to $(-1)^{s}\alpha+t\pi$ where $s,t\in\{0,1\}$. This feed-forward mechanism is used to address the non-determinism of quantum measurement outcomes.
MBQC has the same computation power with the circuit model in the sense that they are both universal computing models. There is a straightforward translation (translation, ) from a circuit in the universal gate set $\{J(\alpha),\mathrm{CZ}\}$ into a measurement pattern on a graph state, where
$J(\alpha)=\begin{bmatrix}1&\mathrm{e}^{\mathrm{i}\alpha}\\ 1&-\mathrm{e}^{\mathrm{i}\alpha}\end{bmatrix}$
For example, Fig. 3 shows the translation from  3(a) to  3(b), each vertex in (b) representing a qubit, with ‘in’ and ‘out’ denoting their roles of being input or output qubits. It can be seen that the gates $J(\alpha)$, $J(\beta)$$,J(\gamma)$ are translated to equatorial measurements with corresponding angles, i.e., $E(\alpha)$, $E(\beta)$, $E(\gamma)$, while the CZ gates are translated to edges of the graph states. This process can be rigorously described in ZX-calculus and optimized by available tools such as PyZX (kissinger2019pyzx, ).
(a)                              (b)
Figure 3. Translation from a circuit (a) to a measurement pattern on a graph state (b).
2.2. Photonic Platform
The weak interaction between photons, despite ensuring low cross-talk between photonic qubits, also poses significant challenges for realizing multi-qubit gates in the circuit model. Therefore, as a computing model that only requires measurements, MBQC (con-rev, ) emerges as highly suitable for photonic quantum computing. Besides the experimental demonstration of small-scale photonic MBQC (mbqc-grover, ; mbqc-dj, ; mbqc-simon, ), the photonic platform is rapidly scaling up with integrated waveguides and optical chips  (con-rev-217, ; con-rev-204, ; ferreira2022deterministic, ; con-rev-247, ; con-rev-252, ; interleaving, )
Practical photonic hardware scales up by creating small resource states, e.g., 4-qubit, 6-qubit graph states, and connecting them through fusions (fbqc, ). In particular, identical resource states are periodically generated by an array of resource state generators (RSGs) every cycle, with those generated in the same RSG cycle forming a 2D resource state layer (RSL). Along with the time dimension, this creates a 3D array of resource states in the space-time.
The resource states can then be merged probabilistically into larger graph states through (type II (linear_optics, )) fusions, which can be regarded as concurrent measurements of $X\otimes Z$ and $Z\otimes X$, on two photonic qubits from different resource states. Resource states on the same RSL can fuse with their neighbors by a spatial routing of photons, while resource states generated by the same RSG but on different RSLs can fuse with each other by a temporal routing that controls the arrival times of photons at measurement devices.
(a)                              (b)
Figure 4. Form a large RSL from multiple small RSLs.
With the advanced integrated silicon photonics, hardware components described above can operate on the scales characterized by GHz clock rates (GHz_1, ; GHz_2, ; GHz_3, ), potentially leading to a time scale $\sim$ 1 ns for RSG cycles (interleaving, ). Spatial routing can be adjusted in every RSG cycle with switches, while temporal routing can be achieved by temporarily storing photonic qubits in a high-capacity quantum memory known as delay lines, realized by optical fiber technology. With a low transmission loss rate of $<5\%$ per km (fiber, ), photons can have a lifetime of around 5000 RSG cycles in the delay lines. Moreover, the size of RSL is not completely constrained by the number of RSGs, but can be extended by leveraging the tradeoff between spatial and temporal fusions (interleaving, ). For example, the large RSL depicted in Fig. 4(b) can be formed by fusing the edges of several small RSLs as depicted in Fig. 4(a), resembling folding a paper by twice (4(b)$\rightarrow$ 4(a)). With a photon lifetime around 5000 RSG cycles, this allows for an extension of RSL size by up to 5000 times.
However, as the key operation for merging resource states, fusions are intrinsically probabilistic. By allowing two ancilla photons, their success probability can be practically boosted to 75% (0.78, ; 0.95with30, ). While no conceptual limit has been found yet, so far the maximum known success probability attainable using linear optics is 78% by injecting 8 ancilla single photons (0.78, ). Reaching a higher success probability not only requires a larger number of ancilla photons but could also require the ancilla photons to be entangled. For example, it would take 30 entangled ancilla photons to reach a success probability over 95% (0.95with30, ).
3. Motivation and Overview
In this section, we present a motivating example to demonstrate that a straightforward adaption of OneQ is insufficient to yield a scalable compiler in the presence of fusion failures. Then we provide an overview of our innovative randomness-aware compiler designed to effectively overcome these challenges.
3.1. Motivating Example
@@
Fig. 5(c) illustrates OneQ’s strategy with an example of generating a graph state in Fig. 5(a) from a single RSL depicted in Fig. 5(b), with the strategy represented more compactly by Fig. 5(d). Since the resource states have a star-like tree structure, we refer to the qubits of degree 1 as leaf qubits, and those of degree $>1$ as root qubits. Light blue lines in Fig. 5(d) denote leaf-leaf fusions (i.e., fusions between two leaf qubits) of the resource states (yellow circles), while dark blue lines denote root-leaf fusions (i.e., fusions between a root and a leaf qubit).
To handle real-time randomness, a straightforward adaption is to introduce a retry mechanism. For example, the strategy in Fig. 5(c) can result in a dynamic implementation in Fig. 5(e) according to the fusion successes and failures (green and red ellipses), with these fusions performed sequentially from t1 to t11. If a fusion such as t3 fails, we retry the fusion using another two qubits at t4, and the same approach is applied to t6 and t7. This allows us to successfully generate the graph state in Fig. 5(a).
However, it is worth noting that some fatal failures may necessitate the retry of the entire compilation. For example, in Fig. 5(f), the triangular structure ABC is successfully generated from t1 to t4, but subsequent failures at t5 and t6 deplete the qubits in ABC, only leaving the isolated qubit B. In Fig. 5(g), a 5-qubit linear graph state forms from t1 to t9, which provides the potential for generating the triangle ABC if the fusion at t10 succeeds in fusing the two qubits at the line ends. Unfortunately, this fusion fails and consumes the last neighboring resource state of qubit A (except C), leaving A with no chance to fuse with other qubits.
@@
From this example, we can find some critical issues of this dynamic retry mechanism. First, adapting to prior fusion outcomes necessitates a sequential execution of fusions. This considerably extends the processing time for each RSL, resulting in a time inefficiency as subsequent RSLs must wait for the completion of the current one. Second, since the decision-making process for responding to prior fusions occurs in real time, this extended processing time could exceed the limited lifetime of photons, especially for large RSLs. This would result in substantial photon loss, compromising the overall fidelity as computing scales up. Third, the frequent retries in real-time implementation lead to significant deviations from the planned strategy in Fig. 5(c). This undermines the benefits of the proactive planning, eroding the efficiency achieved by the mapping strategy of OneQ.
3.2. Framework Overview
Tolerating randomness in the compilation while maintaining efficiency presents a significant challenge. To address this, we propose an innovative framework that achieves scalability and efficiency simultaneously through a synergy of online and offline passes. The online pass prioritizes the real-time scalability by maximizing the concurrency among fusions and the parallelism of the associated path searching. The offline pass focuses on the efficient deployment of high-level program graph states onto the randomness-eliminated computing resource guaranteed by the online pass. The bridge between the online and offline passes is established through an intermediate software layer positioned between the low-level physical layer and the high-level program layer. This is achieved through a novel FlexLattice IR, along with an instruction set supported by the online pass and fulfilling the requirements of the offline pass.
To provide a concise overview, we exemplify the compilation flow by compiling a simple program graph state in Fig. 6(a) onto the hardware in Fig. 6(b), which is 3 layers of that in Fig. 5(b). Indeed, while Fig. 5(b) depicts only a single RSL, the incorporation of additional layers is both allowed and necessary for larger graph states. Steps (b)$\rightarrow$(d)$\rightarrow$(c) demonstrate the online pass, while step (a)$\rightarrow$(c) illustrates the offline pass. In the online pass, fusions are conducted concurrently in a predetermined pattern (Fig. 6(d)) without individual retries of the failed ones, which eliminates the necessity for sequential operations. In this simple example, the resource states would result in a $3\times 3$ lattice if all fusions succeed, since the 3 resource states on the same locations of different layers would form a 4-degree star-like graph state (as depicted in the legend), while these star-like graph states would be joined into a lattice. In the presence of fusion failures, the resulting physical graph state becomes a subgraph of the $3\times 3$ lattice, which is then reshaped to a smaller lattice (Fig. 5(c)). The target structure of the reshaping is program-agnostic, with its simple and regular structure facilitating the enhancement of real-time efficiency. When the fusion success probability exceeds the percolation threshold, this reshaping process attains near-deterministic success as the RSL size increases. This eliminates the necessity for repetitive retries of the entire compilation. With this near-determinism, the offline pass can be employed to improve the efficiency by mapping the program graph state compactly onto the reshaped lattice (bold blue lines in Fig. 5(c)).